# Deep Reinforcement Learning for Infinite Horizon Mean Field Problems in Continuous Spaces

### v1

@@ Abstract
Keywords: Actor-critic, Linear-quadratic control, Mean field game, Mean field control, Mixed mean field control game, Score matching, Reinforcement learning, Timescales.
@@ 1 Introduction
Although traditional numerical methods for solving MFG and MFC problems have proceeded along two avenues, solving a pair of coupled partial differential equations (PDE) [Carmona and Laurière, 2021] or a forward-backward system of stochastic differential equations (FBSDE) [Angiuli et al., 2019], there has been growing interest in solving mean field problems in a model-free way [Angiuli et al., 2022b, Angiuli et al., 2023c, Guo et al., 2019, Perrin et al., 2021, Laurière et al., 2022, Carmona et al., 2021]. With this in mind, we turn to reinforcement learning (RL), an area of machine learning that trains an agent to make optimal decisions through interactions with a “black box” environment. RL can be employed to solve complex problems, such as those found in finance, traffic control, and energy management, in a model-free manner. A key feature of RL is its ability to learn from trial-and-error experiences, refining decision-making policies to maximize cumulative rewards. Temporal difference (TD) methods [Sutton, 1988] are a class of RL algorithms that are particularly well-suited for this purpose. They estimate value functions by updating estimates based on differences between successive time steps, combining the benefits of both dynamic programming and Monte Carlo approaches for efficient learning without requiring a complete model of the environment. For a comprehensive overview of the foundations and numerous families of RL strategies, consult [Sutton and Barto, 2018]. Actor-critic (AC) algorithms—the modern incarnations of which were introduced in [Degris et al., 2012]—are a popular subclass of TD methods where separate components, the actor and the critic, are used to update estimates of both a policy and a value function. The actor is responsible for selecting actions based on the current policy, while the critic evaluates the chosen actions and provides feedback to update the policy. By combining the strengths of both policy- and value-based approaches, AC algorithms achieve more stable and efficient learning.
@@
The mean field term itself poses an interesting problem regarding how to numerically store and update a probability measure on a continuous space in an efficient manner. Some authors have chosen to discretize the continuous space, which leads to a vectorized representation as in [Angiuli et al., 2022b, Angiuli et al., 2023c], while others have looked towards deep learning and deep generative models [Perrin et al., 2021]. We extend the latter avenue by considering a method of distributional learning known as score-matching [Hyvärinen, 2005] in which a probability distribution is represented by the gradient of its log density, i.e., its score function. A parametric representation of the score function is updated using samples from the underlying distribution and allows us to compute new samples from the distribution using a discrete version of Langevin dynamics. We explain how to modify the score-matching procedure for our online regime in Section 4.1.
Building off of the work of [Angiuli et al., 2022b, Angiuli et al., 2023c], in which the authors adapt tabular Q-learning [Watkins, 1989] to solve discrete-space MFG and MFC problems, this paper introduces an AC algorithm in the style of advantage actor-critic [Mnih et al., 2016] for solving continuous-space mean field problems in a unified manner. That is to say, for a given mean field problem, we use the same algorithm to solve for both the MFG and MFC solutions simply by adjusting the relative learning rates of the parametric representations of the actor, critic, and mean field distribution. Our method combines the mean field with the actor-critic paradigm by concurrently learning the score function of the mean field distribution along with the optimal control, which we derive from the policy learned by the actor.
@@
2 Infinite Horizon Mean Field Problems
@@
In both cases, the mathematical setting is a filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ satisfying the usual conditions which supports an $m$-dimensional Brownian motion $(W_{t})_{t\geq 0}$. The measurable function $f:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}$ is known as the running cost, and $\beta>0$ is a discount factor. For the state dynamics we have drift $b:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}^{d}$ and volatility $\sigma:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}^{d\times m}$.
@@ 3.1 Temporal Difference Methods
Temporal difference methods center around iteratively updating an approximation $V$ to $v_{\pi}$ in order to sufficiently minimizes the TD error $\delta_{n}$,
@@ 3.2 Actor-Critic Algorithms
Actor-critic algorithms form a subset of TD methods in which explicit representations of the policy (the actor) and the value function (the critic) are stored. Often, the representation of the policy is a parametric family of density functions in which the parameters are the outputs of another parametric family of functions, e.g., linear functions, polynomials, and neural networks. In the implementation discussed in Section 4, our actor is represented by a feedforward neural network which outputs the mean and standard deviation of a normal distribution. The action $A$ is then sampled according to this density. The benefit of a stochastic policy such as this is that it allows for more exploration of the environment so that the agent does not myopically converge to a suboptimal policy.
@@
$L_{V}(\theta)\coloneqq\left(r_{n+1}+\gamma V_{\theta}(X_{t_{n+1}})-V_{\theta}(X_{t_{n}})\right)^{2}\eqqcolon\delta_{n}^{2}.$ (3.6)
@@
$\nabla_{\theta}L_{V}(\theta)=-2\delta_{n}\nabla_{\theta}V_{\theta}(X_{t_{n}}),$ (3.7)
@@ 4 Unified Mean Field Actor-Critic Algorithm for Infinite Horizon
In this section, we introduce a novel infinite horizon mean field actor-critic (IH-MF-AC) algorithm for solving both MFG and MFC problems in continuous time and continuous space. Although there have been significant strides in recasting the MDP framework for continuous-time using the Hamiltonian of the associated continuous-time control problem as an analog of the Q-function [Jia and Zhou, 2022a, Jia and Zhou, 2022b, Jia and Zhou, 2023, Wang et al., 2020], we instead take the classical approach of first discretizing the continuous-time problem and then applying the MDP strategies discussed in Section 3. As our focus is aimed at identifying the stationary solution of the infinite horizon mean field problems, discretizing time does not meaningfully depart from the original continuous-time problem presented in Section 2; While in our ongoing work [Angiuli et al., 2023b] where we tackle the finite horizon regime, the time-discretization must be treated with more care since the mean field becomes a flow of probability distributions parameterized by time, and the optimal control also becomes time-dependent in this context. In the sequel, we will first recast the mean field setting from Section 2 as a discrete MDP parameterized by the distribution $\mu$ and then lay out the general procedure of the algorithm before addressing the continuous-space representation of $\mu$ via score functions in Section 4.1. Section 4.2 addresses the justification for alternating between the MFG and MFC solutions using the actor, critic, and mean field learning rates.
@@
and the state dynamics in eqs. 2.2 and 2.4 as
@@
This reformulation is directly in correspondence with the MDP setting presented in Section 3—albeit, parameterized by $\mu$. Observe that $r_{t_{n+1}}=-f(X_{t_{n}},\mu,A_{t_{n}})\Delta t$, $\gamma=e^{-\beta\Delta t}$, and the state transition dynamics are given by eq. 4.2.
In the style of the AC method described in Section 3.2, our algorithm maintains and updates a policy $\Pi_{\psi}$ and a value function $V_{\theta}$ which are meant as stand-ins for the control $\hat{\alpha}$ (resp. $\alpha^{*}$) of the MFG (resp. MFC) and the cost functional $J$, respectively. Both are taken to be feedforward neural networks. The third component is the mean field distribution $\mu$, which is updated simultaneously with the actor and critic at each timestep to approximate the law of $X_{t}$. The procedure at the $n^{th}$ step is as follows: the agent is in the state $X_{t_{n}}$ as a result of the dynamics in eq. 4.2 where $\mu$ is replaced with the current estimate of the mean field $\mu_{n-1}$. The value of $X_{t_{n}}$ is then used to update the mean field, yielding a new estimate $\mu_{n}$ (see section Section 4.1 for details). Using the actor’s policy, the agent samples an action $A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$ and executes it in the environment. It receives a reward which, unbeknownst to the agent, is given by $r_{t_{n+1}}=-f(X_{t_{n}},\mu_{n},A_{t_{n}})\Delta t$. The environment places the agent in a new state $X_{t_{n+1}}$ according to eq. 4.2 using the distribution $\mu_{n}$ and the action $A_{t_{n}}$ while $\Pi_{\psi_{n}}$ and $V_{\theta_{n}}$ are updated according to the update rules from Section 3.2. To mimic the infinite horizon regime, we iterate this procedure for a large number of steps until we achieve convergence to the limiting distribution $\hat{\mu}$ (resp. $\mu^{*}$) and the equilibrium (resp. optimal) control $\hat{\alpha}$ (resp. $\alpha^{*}$). The complete pseudocode is presented in Algorithm 1.
@@
0:  Initial distribution $\xi$; number of time steps $N\gg 0$; discrete time step size $\Delta t$; neural network learning rates for actor $\rho_{\Pi}$, critic $\rho_{V}$, and score $\rho_{\Sigma}$; Langevin dynamics step size $\epsilon$.
@@
2:  Sample $X_{t_{0}}\sim\xi$
@@
4:     Compute score loss: $\quad L_{\Sigma}(\varphi_{n})=\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
5:     Update score with SGD: $\quad\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}(\varphi_{n})$
6:     Generate mean field samples $S_{t_{n}}=\left(S_{t_{n}}^{(1)},S_{t_{n}}^{(2)},\dots,S_{t_{n}}^{(k)}\right)$ from $\Sigma_{\varphi_{n+1}}$ using Langevin dynamics (eq. 4.3) with step size $\epsilon$ and compute $\overline{\mu}_{S_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{S_{t_{n}}^{(i)}}$.
7:     Sample action: $\quad A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$
8:     Observe reward from the environment: $\quad r_{n+1}=-f(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},A_{t_{n}})\Delta t$
9:     Observe the next state from the environment:$\quad X_{t_{n+1}}=b(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},A_{t_{n}})\Delta t+\sigma(X_{t_{n}},\overline{\mu}_{S_{t_{n}}},A_{t_{n}})\sqrt{\Delta t}\,z_{n},\qquad z_{n}\sim\mathcal{N}(0,1)$
10:     Compute TD target: $\quad y_{n+1}=r_{n+1}+e^{-\beta\Delta t}V_{\theta_{n}}(X_{t_{n+1}})$
11:     Compute TD error: $\quad\delta_{\theta_{n}}=y_{n+1}-V_{\theta_{n}}(X_{t_{n}})$
12:     Compute critic loss: $\quad L_{V}(\theta_{n})=\delta_{\theta_{n}}^{2}$
13:     Update critic with SGD: $\quad\theta_{n+1}=\theta_{n}-\rho_{V}\nabla_{\theta}L_{V}(\theta_{n})$
14:     Compute actor loss: $\quad L_{\Pi}(\psi_{n})=-\delta_{\theta_{n}}\log\Pi_{\psi_{n}}(A_{t_{n}}\mid X_{t_{n}})$
15:     Update actor with SGD: $\quad\psi_{n+1}=\psi_{n}-\rho_{\Pi}\nabla_{\psi}L_{\Pi}(\psi_{n})$
16:  end for
17:  return  $(\Pi_{\psi_{N}},\Sigma_{\varphi_{N}})$

### v3

@@ Abstract
Keywords: Reinforcement learning, Actor-critic, Mean field game, Mean field control, Mixed mean field control game, Score matching, Timescales, Linear-quadratic problems.
@@ 1 Introduction
Although traditional numerical methods for solving MFG and MFC problems have proceeded along two avenues, solving a pair of coupled partial differential equations (PDE) [Carmona and Laurière, 2023] or a forward-backward system of stochastic differential equations (FBSDE) [Angiuli et al., 2019], there has been growing interest in solving mean field problems in a model-free way [Angiuli et al., 2022b, Angiuli et al., 2023b, Guo et al., 2019, Perrin et al., 2021, Laurière et al., 2022, Carmona et al., 2023]. With this in mind, we turn to reinforcement learning (RL), an area of machine learning that trains an agent to make optimal decisions through interactions with a “black box” environment. RL can be employed to solve complex problems, such as those found in finance, traffic control, and energy management, in a model-free manner. A key feature of RL is its ability to learn from trial-and-error experiences, refining decision-making policies to maximize cumulative rewards. Temporal difference (TD) methods [Sutton, 1988] are a class of RL algorithms that are particularly well-suited for this purpose. They estimate value functions by updating estimates based on differences between successive time steps, combining the benefits of both dynamic programming and Monte Carlo approaches for efficient learning without requiring a complete model of the environment. For a comprehensive overview of the foundations and numerous families of RL strategies, consult [Sutton and Barto, 2018]. Actor-critic (AC) algorithms—the modern incarnations of which were introduced in [Degris et al., 2012]—are a popular subclass of TD methods where separate components, the actor and the critic, are used to update estimates of both a policy and a value function. The actor is responsible for selecting actions based on the current policy, while the critic evaluates the chosen actions and provides feedback to update the policy. By combining the strengths of both policy- and value-based approaches, AC algorithms achieve more stable and efficient learning.
@@
The mean field term itself poses an interesting problem regarding how to numerically store and update a probability measure on a continuous space in an efficient manner. Some authors have chosen to discretize the continuous space, which leads to a vectorized representation as in [Angiuli et al., 2022b, Angiuli et al., 2023b], while others have looked towards deep learning and deep generative models [Perrin et al., 2021]. We extend the latter avenue by considering a method of distributional learning known as score-matching [Hyvärinen, 2005] in which a probability distribution is represented by the gradient of its log density, i.e., its Stein score function. A parametric representation of the score function is updated using samples from the underlying distribution and allows us to compute new samples from the distribution using a discrete version of Langevin dynamics. We explain how to modify the score-matching procedure for our online regime in Section 4.1.
Building off of the work of [Angiuli et al., 2022b, Angiuli et al., 2023b], in which the authors adapt tabular Q-learning [Watkins, 1989] to solve discrete-space MFG and MFC problems, the main contributions of this papers are:
  1. 1.
    We address continuous-state and -action space mean field problems in a unified manner by developing an AC algorithm inspired by advantage actor-critic [Mnih et al., 2016]. This means that, for a given mean field problem, we use the same algorithm to converge to the MFG and MFC solutions simply by adjusting the relative learning rates of the parametric representations of the actor, critic, and mean field distribution. Compared to Q-learning, our proposed algorithm excels in handling continuous or large finite state- and action-spaces, thanks to our neural network implementation. Additionally, it has the advantage of directly learning a policy.
  2. 2.
    To effectively represent the mean field distribution in continuous space, we employ the score function, defined as the gradient of its log density. We parameterize the score function using neural networks and update it via score-matching techniques [Hyvärinen, 2005]. Our method then combines the mean field with the actor-critic paradigm by concurrently learning the score function of the mean field distribution alongside the optimal control, which we derive from the policy learned by the actor.
Note that our approach relies on a randomized policy for the actor while another direction of research utilizes entropy-regularization of the value function as in [Guo et al., 2022, Cui and Koeppl, 2021, Wang et al., 2020]. See also the comment in Section 5.5 describing the relationship between the two approaches. Other papers have concentrated on reinforcement learning for McKean-Vlasov control problems in continuous-time such as in [Wei and Yu, 2023], but our approach treats both MFG and MFC in the same algorithm. We also refer to [Frikha et al., 2023] for learning MFC problems when the policy may depend additionally on the population distribution.
@@
2 Asymptotic Infinite Horizon Mean Field Problems
@@
In both cases, the mathematical setting is a filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ satisfying the usual conditions which supports an $m$-dimensional Brownian motion $(W_{t})_{t\geq 0}$. The measurable function $f:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}$ is known as the running cost, where $\mathcal{P}(\mathbb{R}^{d})$ denotes space of probability measures on $\mathbb{R}^{d}$ and $\mathcal{P}_{s}(\mathbb{R}^{d})$ is the subspace of $\mathcal{P}(\mathbb{R}^{d})$ with finite $s^{th}$-moments. The constant $\beta>0$ is a discount factor, which measures the relative importance or value of future rewards or costs compared to immediate ones. For the state dynamics we have drift $b:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}^{d}$ and volatility $\sigma:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{k}\to\mathbb{R}^{d\times m}$.
@@
It was shown in [Angiuli et al., 2022b] that the asymptotic formulation of the MF problem is the limit of the traditional time-dependent problem as $t\to\infty$ in the following sense: the equilibrium (resp. optimal) control for the time-dependent MFG (resp. MFC) problem converges to the equilibrium (resp. optimal) control for the asymptotic problem as $t\to\infty$.
@@ 3 Reinforcement Learning and Actor-Critic Algorithms
For our continuous-space setting, we randomize the policy according to a Gaussian distribution as explained in Section 3.2.
@@ 3.1 Temporal Difference Methods
Temporal difference methods center around iteratively updating an approximation $V$ to $v_{\pi}$ in order to drive to zero the TD error $\delta_{n}$,
@@ 3.2 Actor-Critic Algorithms
Actor-critic algorithms form a subset of TD methods in which explicit representations of the policy (the actor) and the value function (the critic) are stored. Often, the representation of the policy is a parametric family of density functions in which the parameters are the outputs of another parametric family of functions, e.g., linear functions, polynomials, and neural networks. In the implementation discussed in Section 4, our actor is represented by a feedforward neural network which outputs the mean and standard deviation of a normal distribution. The action $A_{t_{n}}$ is then sampled according to this density. The benefit of a stochastic policy such as this is that it allows for more exploration of the environment so that the agent does not myopically converge to a suboptimal policy.
@@
$L_{V}^{(n)}(\theta)\coloneqq\left(r_{n+1}+\gamma V_{\theta}(X_{t_{n+1}})-V_{\theta}(X_{t_{n}})\right)^{2}\eqqcolon\delta_{n}^{2}.$ (3.6)
@@
$\nabla_{\theta}L_{V}^{(n)}(\theta)=-2\delta_{n}\nabla_{\theta}V_{\theta}(X_{t_{n}}),$ (3.7)
@@ 4 Unified Mean Field Actor-Critic Algorithm for Infinite Horizon
In this section, we introduce a novel infinite horizon mean field actor-critic (IH-MF-AC) algorithm for solving both MFG and MFC problems in continuous-time and continuous-space. Although there have been significant strides in recasting the MDP framework for continuous-time using the Hamiltonian of the associated continuous-time control problem as an analog of the Q-function [Jia and Zhou, 2022a, Jia and Zhou, 2022b, Jia and Zhou, 2023, Wang et al., 2020], we instead take the classical approach of first discretizing the continuous-time problem and then applying the MDP strategies discussed in Section 3. As our focus is aimed at identifying the stationary solution of the infinite horizon mean field problems, discretizing time does not meaningfully depart from the original continuous-time problem presented in Section 2; While in our ongoing work [Angiuli et al., 2024] where we tackle the finite horizon regime, the time-discretization must be treated with more care since the mean field becomes a flow of probability distributions parameterized by time, and the optimal control also becomes time-dependent in this context. In the sequel, we will first recast the mean field setting from Section 2 as a discrete MDP parameterized by the distribution $\mu$ and then lay out the general procedure of the algorithm before addressing the continuous-space representation of $\mu$ via score functions in Section 4.1. Section 4.2 addresses the justification for alternating between the MFG and MFC solutions using the actor, critic, and mean field learning rates.
@@
and the state dynamics in eqs. 2.2 and 2.4 as the Euler-Maruyama approximation
@@
This reformulation is directly in correspondence with the MDP setting presented in Section 3—albeit, parameterized by $\mu$. Observe that $r_{t_{n+1}}=-f(X_{t_{n}},\mu,A_{t_{n}})\Delta t$, $\gamma=e^{-\beta\Delta t}$, and the state transition dynamics are given by eq. 4.2. Note that the model and approximation approach used to derive the reward and next state dynamics are our own choice for simulation purposes but are not prescriptive. The agent sees the environment as a black box and has no knowledge of the details of the problem.
In the style of the AC method described in Section 3.2, our algorithm maintains and updates a policy $\Pi_{\psi}$ and a value function $V_{\theta}$ which are meant as stand-ins for the control $\hat{\alpha}$ (resp. $\alpha^{*}$) of the MFG (resp. MFC) and the cost functional $J$, respectively. Both are taken to be feedforward neural networks. The third component is the mean field distribution $\mu$, which is updated simultaneously with the actor and critic at each timestep to approximate the law of $X_{t}$. The procedure at the $n^{th}$ step is as follows: the agent is in the state $X_{t_{n}}$ as a result of the dynamics in eq. 4.2 where $\mu$ is replaced with the current estimate of the mean field $\mu_{n-1}$. The value of $X_{t_{n}}$ is then used to update the mean field, yielding a new estimate $\mu_{n}$ (see section Section 4.1 for details). Using the actor’s policy, the agent samples an action $A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$ and executes it in the environment. The agent receives a reward which, unbeknownst to the agent, is given by $r_{t_{n+1}}=-f(X_{t_{n}},\mu_{n},A_{t_{n}})\Delta t$ (the value $r_{t_{n+1}}$ is known to the agent but the form of $f$ is not). The environment places the agent in a new state $X_{t_{n+1}}$ according to eq. 4.2 using the distribution $\mu_{n}$ and the action $A_{t_{n}}$. The previous steps encapsulate what is meant by model free in the sense that the agent has no explicit knowledge of the function $f$ nor the state transition dynamics given by $b$ and $\sigma$; it relies only on the immediate information $r_{t_{n+1}}$ and $X_{t_{n+1}}$. Subsequently, $\Pi_{\psi_{n}}$ and $V_{\theta_{n}}$ are updated according to the update rules from Section 3.2. To mimic the infinite horizon regime, we iterate this procedure for a large number of steps until we achieve convergence to the limiting distribution $\hat{\mu}$ (resp. $\mu^{*}$) and the equilibrium (resp. optimal) control $\hat{\alpha}$ (resp. $\alpha^{*}$). The complete pseudocode is presented in Algorithm 1.
@@
0:  Number of time steps $N\gg 0$; discrete time step size $\Delta t$; neural network learning rates for actor $\rho_{\Pi}$, critic $\rho_{V}$, and score $\rho_{\Sigma}$; Langevin dynamics step size $\epsilon$.
@@
2:  Agent receives initial state $X_{t_{0}}$ from the Environment.
@@
4:     Environment computes score loss: $\quad L_{\Sigma}^{(n)}(\varphi_{n})=\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}$
5:     Environment updates score with SGD: $\quad\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}^{(n)}(\varphi_{n})$
6:     Environment generates mean field samples $S_{t_{n}}=\left(S_{t_{n}}^{(1)},S_{t_{n}}^{(2)},\dots,S_{t_{n}}^{(k)}\right)$ from $\Sigma_{\varphi_{n+1}}$ using Langevin dynamics (eq. 4.3) with step size $\epsilon$ and compute $\overline{\mu}_{S_{t_{n}}}\coloneqq\frac{1}{k}\sum_{i=1}^{k}\delta_{S_{t_{n}}^{(i)}}$.
7:     Agent samples action: $\quad A_{t_{n}}\sim\Pi_{\psi_{n}}(\cdot\mid X_{t_{n}})$
8:     Agent observes reward $r_{n+1}$ and next state $X_{t_{n+1}}$ generated from the environment based on its knowledge of $\overline{\mu}_{S_{t_{n}}}$.
9:     Agent computes TD target: $\quad y_{n+1}=r_{n+1}+e^{-\beta\Delta t}V_{\theta_{n}}(X_{t_{n+1}})$
10:     Agent computes TD error: $\quad\delta_{\theta_{n}}=y_{n+1}-V_{\theta_{n}}(X_{t_{n}})$
11:     Agent computes critic loss: $\quad L_{V}^{(n)}(\theta_{n})=\delta_{\theta_{n}}^{2}$
12:     Agent updates critic with SGD: $\quad\theta_{n+1}=\theta_{n}-\rho_{V}\nabla_{\theta}L_{V}^{(n)}(\theta_{n})$
13:     Agent computes actor loss: $\quad L_{\Pi}^{(n)}(\psi_{n})=-\delta_{\theta_{n}}\log\Pi_{\psi_{n}}(A_{t_{n}}\mid X_{t_{n}})$
14:     Agent updates actor with SGD: $\quad\psi_{n+1}=\psi_{n}-\rho_{\Pi}\nabla_{\psi}L_{\Pi}^{(n)}(\psi_{n})$
15:  end for
16:  return  $(\Pi_{\psi_{N}},\Sigma_{\varphi_{N}})$